# Analysis of static and dynamic batching algorithms for graph neural networks

### v2

@@ Abstract
Graph Neural Networks, Batching, Chemistry data, Materials science data
@@ 1 Introduction
Graph neural networks (GNN) models have recently shown great promise in regression and classification tasks, where the input data can be represented as graphs (Kipf & Welling, 2016) (Xu et al., 2018). These methods have been applied to predict molecular and material behavior from nanometer to millimeter scale (Schütt et al., 2018; Jørgensen et al., 2018; Sanchez-Gonzalez et al., 2020; Neumann et al., 2024). These tasks are of upmost importance to society, since they have opened up a research path towards exploring new molecules for drug design and carbon capture (Choudhary et al., 2022) and finding new materials for energy storage and generation. (Schaarschmidt et al., 2022).
Similar to neural network (NN) models, GNN models typically contain millions of parameters and require large training datasets to achieve sufficient predictive power. In order to effectively train GNNs on large datasets, the graph structured data needs to be batched, otherwise each update over the entire dataset takes too long for the model parameters to converge in a reasonable amount of time. In this paper, we examine two different batching techniques for graph networks, static and dynamic batching. The two algorithms differ in that static batching always grabs the same number of graphs whereas dynamic batching adds graphs to the batch conditionally and seeks to ensure that the memory occupied by the batch of graphs is constant.
With the advent of neural architecture searches in different fields, including with graph structured data, it is often the case that researchers train model candidate architectures several thousand times (Zoph, 2016; Gao et al., 2021; Speckhard et al., 2023). End-users are likely to re-train the resulting model and re-tune hyper-parameters on new datasets. The total costs involved with the GNN model search and re-training pose a significant computational, financial and environmental burden (Korolev & Mitrofanov, 2024; Speckhard et al., 2025; Patterson et al., 2021). Therefore, any savings in training time offered by the batching algorithm are significant.
@@
  • We formally introduce the static and dynamic batching algorithms.
  • The different padding schemes determine the number of recompilations for a given dataset and batch size.
  • The performance of the algorithm is dependent on the graph distribution in the dataset, model used, batch size, hardware, and the number of training steps run.
  • Across the experiments, we observe at most a 2.7x speedup in mean time per training step when switching from the slowest algorithm to the fastest.
@@ 2 Related Work
We briefly discuss different batching methods for NNs. The first method is full-batching where the entirety of the dataset is fed through a forward pass of the network. The loss, L, is computed for each of the N datapoints, $(x_{i},y_{i})$, in the dataset D,
$L=\sum_{i=1}^{N}L_{i}(y_{i},x_{i}).$ (1)
An example of the loss, for regression, would be the mean squared error. The gradient of the loss with respect to the model weights, $w$, is used to update the weights. The simplest update equation is gradient descent,
$w^{\prime}=w+\epsilon\nabla_{w}\sum_{i=1}^{N}L_{i},$ (2)
where $\epsilon$ represents the step size. Running the entire dataset through a forward pass of the model and computing the gradients of the weight parameters with respect to the loss is expensive on larger datasets for deep neural networks (DNN) where backpropagation is required (Hastie, 2009). As a result, typically, mini-batches are used for NN training where in each update-step a subset of the data is used to update model weights (Bishop & Nasrabadi, 2006). This leads to a batch gradient descent update equation. In the extreme case, known as stochastic (or on-line learning) gradient descent, the batch size is a single datapoint (Bottou et al., 1991).
@@
These different batching methods have been thoroughly analyzed for NNs in the literature (Bottou & Bousquet, 2007) as a function of the batch size. Computation time as a function of batch size has also been explored on GPUs (Kochura et al., 2019). In (Byrd et al., 2012), large batch sizes were found to typically slow down the convergence of the model parameters. In practice, researchers typically set the batch size as a hyperparameter that is found via cross validation (CV).
@@
For NNs, updating the model weights can take up the bulk of training time through backpropagation (Lister & Stone, 1995). Typically, NNs operate on numeric data (or data that has been transformed into numeric values). When training the NN with a fixed mini-batch size of numeric data, batches have constant shape and memory requirements. Using Jax, this enables highly efficient training, since the the gradient update step can be compiled once at the start of the training. For convolutional neural networks (CNNs), this is not the case, and images of different pixel dimensions are often padded before being fed into the network (Tang et al., 2019). Similarly, for graph neural networks, the graphs in the dataset typically contain a wide variety of number of nodes and edges. To this end, batching algorithms have emerged, which pad batches of graphs to constant shapes. In this way, the gradient update step for GNNs on batches can be compiled on GPU. Two such methods have become popular, static batching and dynamic batching. The static batching methods always collect a fixed number of graphs while dynamic methods add graphs incrementally to a batch until some padding budget/target is reached.
@@
That said, not all models pad their data. M3GNet (Chen & Ong, 2022), a GNN trained for interatomic potentials (IAP) is written in TensorFlow and performs batching in a similar way to NNs. Its batching procedure collects a number of graphs corresponding to the batch size, and concatenates the atoms, bonds, states and indices into larger lists for the batch. However, it does not perform any padding after batching data. The pyTorch GNN (ptgnn) library implements a dynamic batching algorithm that also does not use padding (Allamanis et al., 2022). The algorithm adds graphs until either the batch has the desired batch size (i.e., number of graphs) or some safety limit on the number of nodes in the batch has been reached. This safety limit ensures that the batch will fit into memory.
@@
The pyTorch geometric library (Fey & Lenssen, 2019) implements a similar dynamic batching algorithm. The user specifies whether to use either (but not both) a total number of nodes as the batch cutoff/limit or total number of edges. The algorithm then adds graphs incrementally to the batch until the target number of graphs are in the batch (i.e., the target batch size) or the cutoff has been reached. It also allows the user to skip adding single graphs to the batch that would by themselves exceed the node/edge cutoff.
@@
The Jraph library, which is built on JAX (Bradbury et al., 2018), implements a dynamic batching algorithm (Godwin* et al., 2020). Given a batch size, it performs a one-time estimate of the next largest multiple of 64 to which the sum of the nodes/edges in the batch should be padded. We can think of these estimates as the node/edge padding targets for each batch (i.e., the number of nodes/edges in a batch after padding). This estimation is done by sampling a random subset of the data. It then iteratively adds one graph at a time to the batch and stops if adding another graph would exceed the node/edge padding targets or the maximum number of graphs (i.e., the batch size) is already in the batch.
@@
The Tensorflow GNN library (Ferludin et al., 2023) implements a static and dynamic batching algorithm. The static batching adds a fixed number of graphs to the batch and then pads to a constant padding value. The dynamic batching method, similar to Jraph, estimates a padding budget (they call it a size constraint) for the batch based on a random subset of data, and then adds graphs incrementally to the batch as long as they do not break this budget.
@@
For datasets with larger graphs, different batching methods have emerged. PipeGCN (Wan et al., 2022) breaks larger training graphs into smaller pieces and trains each partition on a different GPU. Node feature information from different partitions is passed between GPUs across a fast cross-GPU link. Batching methods that break training graphs into partitions are referred to in the literature as mini-batching in GNNs (Bajaj et al., 2024) as compared to methods that always deal with the entire graph which are called full graph or full-batch methods. These methods are not dealt with in this paper since the methods are meant for datasets which contain very large graphs, unlike the datasets used in this work.
@@ 3 Preliminaries
We now dive deeper into the static and dynamic batching algorithms. In static batching, the method grabs a fixed number of graphs for each batch. It then checks how many nodes/edges are present in the batch. It then adds a dummy graph and adds dummy edges/nodes to this graph to serve as padding. It pads the number of nodes/edges up to the next power of two to ensure that all batches have the same number of nodes/edges (and graphs), so that the memory of the input to the update function is constant and does not require recompilation. This means that, typically, the number of graphs in the batch is the batch size minus one in order to leave space for the dummy graph in the batch. The fact that a power of two is used for padding is arbitrary. We refer to this algorithm as static-$2^{N}$. We also evaluate static batching with padding to the nearest multiple of 64. We label this algorithm as static-64. A third method is also possible, and implemented in TensorFlow’s GNN (TF-GNN) library, where the algorithm first finds the graph with the most nodes/edges in the graph and multiplies this number by the batch size to get the padding target. We refer to this algorithm as static-constant.
@@
Over the course of training, one might encounter a batch that requires a larger power of two than previously used to contain all of the nodes and edges in the batch. This requires the gradient-update function to recompile for a different, i.e., larger, input size. The static batching algorithm is shown in Algorithm 1. The padding method, pad-power-of-2, called in the algorithm, adds a dummy graph with fake nodes/edges to the list of graphs, so that the sum of the nodes/edges in the list of graphs is a power of two. Pseudocode for the method is found in the Appendix, Section A. The batch method called in the algorithm, concatenates all of the nodes/edges in the list of graphs and creates a single super-graph out of the list of graphs (pseudocode is found in the Appendix).
Algorithm 1 Static batching
  Input: batch size $B_{G}$, training graphs $G$, start point $s$
  $A\leftarrow[]$Accumulated list of graphs}
  $n_{n}\leftarrow 0$Accumulated # of nodes}
  $n_{e}\leftarrow 0$Accumulated # of edges}
  for $i=s$ to $s+B_{G}$ do
     $A\leftarrow\text{$A$.append($G[i]$)}$
[…]
     if $len(A)==(B_{G}-1)$ then
        return batch(pad-nearest-power-2($A$, $n_{n}$, $n_{e}$))
     end if
  end for
@@
The dynamic batching algorithm, as implemented in Jraph and TF-GNN, first grabs a random subset of the data (e.g., 1000 graphs). It then estimates the mean nodes per graph and mean edges per graph. It uses this to create a padding target (or memory budget), which in effect is the maximum number of nodes and edges allowed to be stored in the batch. The edge (node) padding target is created by multiplying the mean number of edges (nodes) by the batch size and then rounding to the nearest multiple of 64. The padding target estimation is shown in Algorithm 2, and is used to determine the padding required for each batch to ensure that each batch uses the same amount of memory.
Algorithm 2 Estimate dynamic batching padding target
  Input: subset size $N$, list of graphs $G$, batch size $B_{G}$
  $n\leftarrow 0$number of nodes in N graphs}
  $e\leftarrow 0$number of edges in N graphs}
  $g\leftarrow G[1:N]$Grab N graphs from training data}
  for $i=1$ to $i=N$ do
     $n$ += $G[i].$num-nodes
     $e$ += $G[i].$num-edges
  end for
  $B_{n}\leftarrow\text{next-multiple-64}(n*B_{G}/N)$padding target}
  $B_{e}\leftarrow\text{next-multiple-64}(e*B_{G}/N)$
  return $B_{n}$, $B_{e}$, $B_{G}$
@@
The dynamic algorithm, given the padding target, appends one graph at a time to the initially empty batch. Before the addition of each graph, it checks if adding the graph would put the batch over the node or edge padding target. If it does, the graph is not added to the batch. Instead, the algorithm adds a dummy graph with fake nodes/edges, the number of which is determined so as to pad the number of nodes/edges to the padding target. It then adds, if necessary, dummy graphs with no nodes/edges to ensure that the number of graphs in the batch is equal to the batch size. A pseudocode representation is shown in Algorithm 3. If, however, during the course of training, a single graph is encountered with more nodes/edges than the padding target, the program terminates. The algorithm must be restarted with a larger node/edge padding target. This issue can be minimized by looking at a larger sample of graphs when estimating the targets. Alternatively, the user can loop through the entire dataset of graphs and find the graph with the largest number of nodes/edges to determine suitable budgets. Note that the algorithm we present here is quite different from the pyTorch-Geometric implementation, whose implementation does not perform a budget estimation (has to be entered by the user) and it does not perform a check on the number of edges, but only the nodes. We find the Jraph/TF-GNN implementation, for which a simplified version is shown in Algorithm 3, to be the more sophisticated, and so it was used for our tests.
@@
        $n_{n}\text{ += }s_{i}[0]$Update node sum}
@@ 4 Datasets
We run profiling experiments on two different datasets. The QM9 dataset (Ramakrishnan et al., 2014) contains chemical properties (e.g., internal energies, molecular orbital energy levels) of small organic molecules. The AFLOW dataset (Curtarolo et al., 2012) is a collection of relevant material properties (e.g., formation energies, band gaps, elastic properties). We choose these datasets for two reasons. First, they have served as benchmark datasets for GNN models (Schütt et al., 2018; Li et al., 2024; Bechtel et al., 2023). Second, they have real world applications: QM9 may help models better perform targeted drug discovery, while AFLOW may help models discover more efficient solar cell semiconductors. For QM9 we target the molecular internal energy and for AFLOW we focus on learning the enthalpy of the crystal.
The datasets are not provided in a graph format. To convert them to graphs, each atom is represented as a node in the graph. For QM9 the edges are added by fully connecting the nodes. For AFLOW, the edges are added with the K-nearest neighbor algorithm based on pairwise distance and setting K equal to 24. This KNN value was chosen from CV studies in the literature (Jørgensen et al., 2018). From the AFLOW dataset, we remove duplicates and keep only calculations that contain both the enthalpy and the band structure, following the procedure in (Bechtel et al., 2023). We visualize the variety in the graph structures in the datasets in Fig. 1. From the figure we can see that the QM9 dataset node distribution appears Gaussian and centered around 17 nodes per graph. AFLOW’s node distribution has long tails despite the mean of the distribution being smaller than QM9. The distribution of edges for AFLOW and QM9 are reflective of the distribution of the nodes, this is because the number of edges are dependent on the number of nodes when using the KNN algorithm or fully connecting the graphs.
Figure 1: Histograms of the number of nodes (top) and edges (bottom) in the AFLOW and QM9 datasets.

### v3

# AI for Accelerated Materials Design

@@ 1 Introduction
Graph neural networks (GNN) models have recently shown great promise in regression and classification tasks, where the input data can be represented as graphs [kipf2016semi, xu2018powerful]. These methods have been applied to predict molecular and material behavior from nanometer to millimeter scale [schutt2018schnet, jorgensen2018neural, sanchez2020learning, neumann2024orb]. These tasks are of upmost importance to society, since they have opened up a research path towards exploring new molecules for drug design and carbon capture [choudhary2022graph] and finding new materials for energy storage and generation [schaarschmidt2022learned].
Similar to neural network (NN) models, GNN models typically contain millions of parameters and require large training datasets to achieve sufficient predictive power. In order to effectively train GNNs on large datasets, the graph structured data needs to be batched, otherwise each update over the entire dataset takes too long for the model parameters to converge in a reasonable amount of time. In this paper, we examine two different batching techniques for graph networks, static and dynamic batching. The two algorithms differ in that static batching always grabs the same number of graphs whereas dynamic batching adds graphs to the batch conditionally and seeks to ensure that the memory occupied by the batch of graphs is constant. For static-batching, we look at three algorithms, the static-64, static-$2^{N}$ and static-constant algorithms.
With the advent of neural architecture searches in different fields, including with graph structured data, it is often the case that researchers train model candidate architectures several thousand times [zoph2016neural, gao2021graph, speckhard2023neural]. End-users are likely to re-train the resulting model and re-tune hyper-parameters on new datasets. The total costs involved with the GNN model search and re-training pose a significant computational, financial and environmental burden [korolev2024carbon, speckhard2025big, patterson2021carbon]. Therefore, any savings in training time offered by the batching algorithm are significant.
@@
  • We formally introduce the static and dynamic batching algorithms and several variants thereof.
@@
  • We find the static-constant algorithm to be the slowest algorithm and recommend not using it unless the goal is to maximize the robustness of the training pipeline. Across our experiments, we observe speedups of up to 12.5x when switching from the static-constant algorithm to another algorithm.
  • Excluding the static-constants batching algorithm, we observe at most a 2.7x speedup in mean time per training step when switching from the slowest algorithm to the fastest.
  • We demonstrate that the performance of the algorithm is dependent on the graph distribution in the dataset, model used, batch size, and the number of training steps run.
  • The static-64 and dynamic algorithms generally outperform the static-$2^{N}$ algorithms.
  • We recommend using the dynamic algorithm, whose performance is similar to the static-64 algorithm, but is faster for fewer number of training steps.
@@ 2 Related Work
Several different batching methods, i.e., full-batching, mini-batching [bishop2006pattern], stochastic learning [bottou1991stochastic], have been thoroughly analyzed for NNs in the literature [bottou2007tradeoffs]. For instance, NN learning performance when using mini-batching has been examined as a function of the batch size. Computation time as a function of batch size has also been explored on GPUs [kochura2019batch]. In Ref. [byrd2012sample], large batch sizes were found to typically slow down the convergence of the model parameters. In practice, researchers typically set the batch size as a hyperparameter that is found via cross validation (CV).
@@
For NNs, updating the model weights can take up the bulk of training time through backpropagation [lister1995empirical]. Typically, NNs operate on numeric data (or data that has been transformed into numeric values). When training the NN with a fixed mini-batch size of numeric data, batches have constant shape and memory requirements. Using JAX, this enables highly efficient training, since the gradient update step can be compiled once at the start of the training. For convolutional neural networks (CNNs), this is not the case, and images of different pixel dimensions are often padded before being fed into the network [tang2019impact]. Similarly, for graph neural networks, the graphs in the dataset typically contain a wide variety of number of nodes and edges. To this end, batching algorithms have emerged, which pad batches of graphs to constant shapes. In this way, the gradient update step for GNNs on batches can be compiled on GPU. Two such methods have become popular, static batching and dynamic batching. The static batching methods always collect a fixed number of graphs while dynamic methods add graphs incrementally to a batch until some padding budget/target is reached.
@@
That said, not all GNN models and libraries pad their data. M3GNet [chen2022universal], a GNN trained for interatomic potentials (IAP) is written in TensorFlow and performs batching in a similar way to NNs. Its batching procedure collects a number of graphs corresponding to the batch size, and concatenates the atoms, bonds, states and indices into larger lists for the batch. However, it does not perform any padding after batching data. The pyTorch GNN (ptgnn) library implements a dynamic batching algorithm that also does not use padding [ptgnn2022microsoft]. The algorithm adds graphs until either the batch has the desired batch size (i.e., number of graphs) or some safety limit on the number of nodes in the batch has been reached. This safety limit ensures that the batch will fit into memory.
@@
The pyTorch geometric library [fey2019fast] implements a similar dynamic batching algorithm. The user specifies whether to use either (but not both) a total number of nodes as the batch cutoff/limit or total number of edges. The algorithm then adds graphs incrementally to the batch until the target number of graphs are in the batch (i.e., the target batch size) or the cutoff has been reached. It also allows the user to skip adding single graphs to the batch that would by themselves exceed the node/edge cutoff.
@@
The Jraph library, which is built on JAX [jax2018github], implements a dynamic batching algorithm [jraph2020github]. Given a batch size, it performs a one-time estimate of the next largest multiple of 64 to which the sum of the nodes/edges in the batch should be padded. We can think of these estimates as the node/edge padding targets for each batch (i.e., the number of nodes/edges in a batch after padding). This estimation is done by sampling a random subset of the data. It then iteratively adds one graph at a time to the batch and stops if adding another graph would exceed the node/edge padding targets or the maximum number of graphs (i.e., the batch size) is already in the batch.
@@
The Tensorflow GNN library [ferludin2023tfgnn] implements a static and dynamic batching algorithm. The static batching adds a fixed number of graphs to the batch and then pads to a constant padding value. The dynamic batching method, similar to Jraph, estimates a padding budget (they call it a size constraint) for the batch based on a random subset of data, and then adds graphs incrementally to the batch as long as they do not break this budget.
@@ 3 Preliminaries
Static-constant batching. In static batching, the method grabs a fixed number of graphs for each batch. It then checks how many nodes/edges are present in the batch. It then adds a dummy graph and adds dummy edges/nodes to this graph to serve as padding. Typically, the number of graphs included in the batch is the batch size minus one in order to leave space for the dummy graph in the batch. The algorithm pads the number of nodes/edges up to the next power of two to try to ensure that most batches have a similar number of nodes/edges, and thereby a similar memory footprints. In the static-constant algorithm, implemented in TensorFlow’s GNN (TF-GNN) library, the algorithm first finds the graph with the most nodes/edges in the dataset and multiplies this number by the batch size to get the padding target. The padding target dictates how many nodes/edges are expected in each batch. For example, for batch size 32, the algorithm grabs 31 graphs, and then adds a dummy graph with enough dummy nodes/edges so that the batch has the same nodes/edges as the padding target. In this way, each static-constant batch occupies the same memory footprint. This means that the gradient-update step, which updates the model parameters, can always expect a batch size of a constant memory input, and can be compiled for this input size. The downside to this approach, is that for some datasets and batch sizes the target is extremely large since it is based off the largest graph in the dataset.
@@
Static-$2^{N}$ and static-64 algorithms. In the static-$2^{N}$ algorithm, which we introduce here, the algorithm grabs batch size minus one number of graphs into a batch, and then pads the batch’s total number of nodes/edges to the next largest power of two. Over the course of training, one might encounter a batch that requires a larger power of two than previously used in order to contain all of the nodes and edges in the batch. This requires the gradient-update function to recompile for a different, i.e., larger, input size. The static-$2^{N}$ algorithm is shown in the appendix Algorithm 4. Similarly, in the static-64 algorithm, which we also introduce here, the total number of nodes/edges in the batch are padded to the next largest multiple of 64.
@@
Dynamic batching. The dynamic batching algorithm, as implemented in Jraph and TF-GNN, first grabs a random subset of the data (e.g., 1000 graphs). It then estimates the mean nodes per graph and mean edges per graph. It uses this to create a padding target (or memory budget), which in effect is the maximum number of nodes and edges allowed to be stored in the batch. The edge (node) padding target is created by multiplying the mean number of edges (nodes) by the batch size and then rounding to the next largest multiple of 64. The algorithm to estimate the padding target is shown in the appendix in Algorithm 5. It is used to determine the padding required for each batch to ensure that each batch uses the same amount of memory.
@@
The dynamic algorithm, given the padding target, appends one graph at a time to the initially empty batch. Before the addition of each graph, it checks if adding the graph would put the batch over the node or edge padding target. If it does, the graph is not added to the batch. Instead, the algorithm adds a dummy graph with fake nodes/edges, the number of which is determined so as to pad the number of nodes/edges to the padding target. It then adds, if necessary, dummy graphs with no nodes/edges to ensure that the number of graphs in the batch is equal to the batch size. A pseudocode representation is shown in Algorithm 1. The batch method called in the algorithm, concatenates all of the nodes/edges in the list of graphs and creates a single super-graph out of the list of graphs (Pseudocode for the method is found in the Appendix, Section B).
@@
   $n$ += $_{n}\text{ += }s_{i}[0$] {Update node sum}
@@
If, however, during the course of training, a single graph is encountered with more nodes/edges than the padding target, the program terminates. The algorithm must be restarted with a larger node/edge padding target. This issue can be minimized by looking at a larger sample of graphs when estimating the targets. Alternatively, the user can loop through the entire dataset of graphs and find the graph with the largest number of nodes/edges to determine suitable budgets. Note that the algorithm we present here is quite different from the pyTorch-Geometric implementation, whose implementation does not perform a budget estimation (has to be entered by the user) and it does not perform a check on the number of edges, but only the nodes. We find the Jraph/TF-GNN implementation, for which a simplified version is shown in Algorithm 1, to be the more sophisticated method, and so it was used for our tests.
@@ 4 Datasets
We run profiling experiments on two different datasets. The QM9 dataset [ramakrishnan2014quantum] contains chemical properties (e.g., internal energies, molecular orbital energy levels) of small organic molecules. The AFLOW dataset [curtarolo2012aflow] is a collection of relevant material properties (e.g., formation energies, band gaps, elastic properties). We choose these datasets for two reasons. First, they have served as benchmark datasets for GNN models [schutt2018schnet, li2024quantum, bechtel2023band]. Second, they have real world applications. For instance, learning on the QM9 dataset may help models better perform targeted drug discovery, while AFLOW data may help models discover more efficient solar cell semiconductors. For QM9 we target the molecular internal energy and for AFLOW we focus on learning the enthalpy of the crystal.
The datasets are not provided in a graph format. To convert them to graphs, each atom is represented as a node in the graph. For QM9 the edges are added by fully connecting the nodes. For AFLOW, the edges are added with the K-nearest neighbor algorithm based on pairwise distance and setting K equal to 24. This KNN value was chosen from CVstudies studies in the literature [jorgensen2018neural]. From the AFLOW dataset, we remove duplicates and keep only calculations that contain both the enthalpy and the band structure, following the procedure in [bechtel2023band]. We visualize the variety in the graph structures in the datasets in Fig. 1. From the figure we can see that the QM9 dataset node distribution appears Gaussian and centered around 17 nodes per graph. AFLOW’s node distribution has long tails despite the mean of the distribution being smaller than QM9’s. The distribution of edges for AFLOW and QM9 are reflective of the distribution of the nodes, this is because the number of edges are dependent on the number of nodes when using the KNN algorithm or fully connecting the graphs.
Figure 1: Left: Histograms of the number of nodes (top) and edges (bottom) in the AFLOW and QM9 datasets. Right: Histograms of the number of nodes (top) and edges (bottom) in a batch of size 32 before padding for the static-$2^{N}$ and dynamic batching algorithms running on the QM9 dataset.